# SVEMnet: An R package for Self-Validated Elastic-Net Ensembles and Multi-Response Optimization in Small-Sample Mixture–Process Experiments

### v1

@@ Appendix B Supplementary multiple-comparison tables
For all Tukey–Kramer analyses we fixed the expansion at the correctly specified quadratic order (Order 2) and fit linear mixed models with Setting as the primary factor of interest, $n_{\text{total}}$, Target_R2, and their interaction as categorical fixed effects, and simulation run ID as a random intercept. Pairwise differences between settings were assessed using Tukey–Kramer honestly significant difference comparisons of the model-based least-squares means, controlling the familywise error rate at $\alpha=0.05$. Settings that do not share a letter differ significantly, and smaller least-squares means correspond to better performance (lower log-NRMSE or log-loss). Here relaxTRUE/FALSE indicates whether relaxed base learners are used, lasso/default indicates the pure lasso case ($\alpha=1$) versus the mixed-$\alpha$ grid $\alpha\in\{0.5,1\}$, and in Table LABEL:tab:tukey-hsd the suffix dbTRUE/FALSE indicates whether the post-hoc debiasing step is applied. In Table LABEL:tab:tukey-binomial all configurations use dbFALSE.
@@
During the preparation of this manuscript and the accompanying software, the author used OpenAI GPT models as assistive tools. Under the author’s direction, these models were used (i) to revise and proofread prose and (ii) to draft the majority of the initial SVEMnet R code and associated documentation. All AI-generated text and code were reviewed, edited, and validated by the author before inclusion. All algorithmic design choices, statistical methodology, simulation design, and final implementation decisions were made by the author, who takes full responsibility for the content of the manuscript, the software, and all reported results.
@@
JMP add-in. An optional JMP add-in, “Relaxed Elastic Net and Lasso SVEM using R” (https://marketplace.jmp.com/appdetails/Relaxed+Elastic+Net+and+Lasso+SVEM+using+R), is available via the JMP Marketplace. The add-in lets JMP users install SVEMnet in R, call the package for a saved Fit Model table script with Gaussian responses, and obtain prediction and standard error formula columns. It provides a convenience front-end to SVEMnet but is not required to reproduce the results in this paper.

### v2

@@ Appendix B Supplementary multiple-comparison tables
For all Tukey–Kramer analyses we fixed the expansion at the correctly specified quadratic order (Order 2) and fit linear mixed models in JMP Pro 19 with Setting as the primary factor of interest, $n_{\text{total}}$, Target_R2, and their interaction as categorical fixed effects, and simulation run ID as a random intercept. Pairwise differences between settings were assessed using Tukey–Kramer honestly significant difference comparisons of the model-based least-squares means, controlling the familywise error rate at $\alpha=0.05$. Settings that do not share a letter differ significantly, and smaller least-squares means correspond to better performance (lower log-NRMSE or log-loss). Here relaxTRUE/FALSE indicates whether relaxed base learners are used, lasso/default indicates the pure lasso case ($\alpha=1$) versus the mixed-$\alpha$ grid $\alpha\in\{0.5,1\}$, and in Table LABEL:tab:tukey-hsd the suffix dbTRUE/FALSE indicates whether the post-hoc debiasing step is applied. In Table LABEL:tab:tukey-binomial all configurations use dbFALSE.
@@
During the preparation of this manuscript and the accompanying software, the author used OpenAI GPT models as assistive tools. Under the author’s direction, these models were used (i) to revise and proofread prose and (ii) to help draft the SVEMnet R code and associated documentation. All AI-generated text and code were reviewed, edited, and validated by the author before inclusion. All algorithmic design choices, statistical methodology, simulation design, and final implementation decisions were made by the author, who takes full responsibility for the content of the manuscript, the software, and all reported results.
@@
JMP add-in. An optional JMP add-in, “Relaxed Elastic Net and Lasso SVEM using R” (https://marketplace.jmp.com/appdetails/Relaxed+Elastic+Net+and+Lasso+SVEM+using+R), is available via the JMP Marketplace. The add-in lets JMP users call SVEMnet for a saved Fit Model table script with Gaussian responses and obtain prediction and standard error formula columns.
Independently tested by
Daniel Fortune, Hexion, Inc., Katy, Texas, USA, installed SVEMnet v3.1.4 from a local .zip file in R 4.5.2 and ran the supplemental LNP_example_SVEMnet.R code shown in Section 3 (and available for download from the repository [40]). The $p$-values printed to the console match those reported in Table 1. The CSV file of candidates saved successfully and contained the same entries (where candidate_type is equal to best) as shown in Table 2 and Table 3.
Acknowledgements
The author thanks Daniel Fortune for verifying installation and the core workflow.